The position of a particle moving along the x-axis is determined by the

equation

d^2x/dt^2+8x=20cos2t

. If the particle starts from rest at x = 0,



determine

(i) The position x as a function of time

(ii) The amplitude, period and frequency of oscillation after a long time has elapsed
1
Expert's answer
2020-11-17T13:03:10-0500
"\frac{d²x}{dt²} + 8x = 20\cos2t."

The solution



"x=A\cos2t+B\sin2t."


We obtain



"-4A\cos2t-4B\sin2t - 8A\sin2t+8B\cos2t""+ 8A\cos2t+8B\sin2t = 20\cos2t."

Hence



"4A+8B=20,\quad -8A+4B=0.""B=2A, \; A=1.""x=\cos2t+2\sin2t=\sqrt{5}\cos(2t+\arctan2).""x=\cos2t+2\sin2t=\sqrt{5}\cos(2t+\arctan2)."

Therefore:

amplitude "a=\sqrt{5};"

"a=\sqrt{5};"

period "T=\frac{2\pi}{\omega}=\pi;"

frequency "f=1/T=1/\pi."